OpenStudy (anonymous):

solve the equation 2cos^2θ-3cosθ+1=0, 0≤θ<0 I believe it has to be in radians. If you could rather help me to the answer, that would be amazing!

OpenStudy (anonymous):

\[2\cos ^{2}\theta-3\cos \theta+1=0\]

OpenStudy (anonymous):

\[0\le \theta <0\]

OpenStudy (anonymous):

can you provide more info like answer choices and what are we doing with this problem explain a little

OpenStudy (anonymous):

There are no answer choices

OpenStudy (anonymous):

that is alright, so what is the problem asking for us to do?

OpenStudy (anonymous):

just to solve for theta

OpenStudy (anonymous):

let me ask you how would you solve for x? what steps must taken? I know the answer I want to see if you know I am trying to help here

OpenStudy (anonymous):

would it be simplified to \[\cos \theta(\cos \theta+1)=2/3\]?

OpenStudy (anonymous):

yes now what?

OpenStudy (anonymous):

how do you get theda by itself

OpenStudy (anonymous):

This is where I become lost.

OpenStudy (anonymous):

wait, would you use arccos to 2/3?

OpenStudy (anonymous):

\[\cos \theta+1=\cos^{-1} (2/3)\]

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

cos th+1 and sub -1

OpenStudy (anonymous):

than divide by cos so you have theda =

OpenStudy (anonymous):

so \[\theta=\cos^{-1} (2/3)-1/\]

OpenStudy (anonymous):

\[\theta=\cos^{-1} (2/3)-1/\cos\]

OpenStudy (anonymous):

yes now it is not done what can be to 2/3-1?

OpenStudy (anonymous):

-1/3?

OpenStudy (anonymous):

yes so what is the final answer now

OpenStudy (anonymous):

Does cos and cos^-1 cancel out leaving only -1/3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Dang.. I do not know what to do after this.

OpenStudy (anonymous):

you are done

OpenStudy (anonymous):

\[\theta=\cos^{-1} (-1/3)/\cos\] Nothing need to be placed next to cos?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok. Thank you :)

OpenStudy (anonymous):

did you get what it was asking for my itself? if yes than you are finished

OpenStudy (anonymous):

Yes :)